# Pushing the limits of eccentricity in planet-disc interactions

### v1

@@ 4.3 Integrated torques
As anticipated by the agreement observed in the radial torque density curves examined in Figs. 3 and 4, here we also find close accordance in the integrated torque across all eccentricities and disc profiles. Notably, the simulations confirm the qualitative dependence on $e$ as previously discussed in detail by FR25. Namely, for the steeper $p=1.5$ runs there is an initial enhancement in $T$ as one approaches and exceeds the transonic value of $e=h_{\rm p}$. The simulations are specifically targeted to capture this peak torque around $e=0.06$, finding almost perfect overlap with the linear predictions. Meanwhile we see that for the shallower surface density profile $p=0.5$, this enhancement is less pronounced and the $T$ remains relatively flat up towards $e=h_{\rm p}$. Beyond this transonic point, we observe a clear transition in behaviour as the net torque undergoes a sharp reversal becoming highly negative and minimised when $e\sim 2h_{\rm p}$. This corresponds to the value of eccentricity for which the planet moves supersonically with respect to the gas at all orbital phases.
@@
It was at this value of $e=0.12$ that FR25 truncated their study due to uncertainty in the modal convergence for higher eccentricities. However, here we push much further and extend our simulations and linear theory up towards $e=0.7$. The net torque remains negative but slowly decreases in absolute value as $e$ increases. Fitting a power law between $e=0.18$ and $e=0.7$ to our linear theory results, yields the best fitting exponents $(-1.1,-1.1,-1.0)$ for the disc profiles $(p,q)=$ $(1.5,1.0)$, $(1.5,0.0)$ and $(0.5,0.0)$ respectively. This scaling is close to the heuristically motivated fitting functions presented in Ida et al. (2020), which suggest that $T\propto e^{-1}$ in the highly supersonic limit. Once again, the broad trends are in good agreement between the methodologies. However, for the highest values of $e$ we begin to see larger discrepancies emerge as the linear theory underestimates the torque magnitude found in the simulations. To test this further, we have performed a convergence study in appendix A which demonstrates that our linear framework and simulation torque diagnostics do not change significantly upon including more modes or doubling the fiducial resolution. This suggests that the differences arise due to more intrinsic difficulties. Perhaps the high value of $e$ fundamentally precludes convergence of our modal scheme. Furthermore, the very oscillatory nature of the wave contributions for high order modes becomes mutually challenging for the linear theory ODE solvers and the fully numerical hydro scheme. Whilst the large radial planetary excursions potentially lead to more troublesome interaction with the boundaries, we show in appendix A.3 that the differences are primarily sourced between apocentre and pericentre. The intricate torque structures in this interval might point towards subtle nonlinear effects that affect the disc response in this region. For example, in section 4.5 we show that nonlinear wave steepening can deposit angular momentum into the disc, which will modify the background density structure slightly, potentially altering the interaction with the planet.
@@ 4.4 Orbital evolution
In Figs. 7 and 8 we examine this orbital element interplay for $\tau_{a}^{-1}$ and $\tau_{e}^{-1}$ respectively, as a function of $e$. Once again, the three rows in each figure correspond to the different disc profiles. Meanwhile the yellow markers and connective lines correspond to the linear theory predictions, whilst the blue stars mark the simulation results. The results are plotted in units of the characteristic migration rate $\tau_{0}^{-1}\equiv F_{J,0}/(M_{\rm p}\Omega_{\rm k,p}a_{\rm p}^{2})$.
@@
Now turning our attention to Fig. 8, we examine the behaviour of the characteristic eccentricity damping rate $\tau_{e}^{-1}$. Firstly, note the different y-axis scale which is roughly two orders of magnitudes larger compared with Fig. 7, consistent with our expectation that the $\tau_{e}^{-1}$ is much faster than $\tau_{a}^{-1}$ (as found previously in FR25). Once again we see the simulations and linear results are in very good agreement. The characteristic eccentricity damping exhibits a sharp rise which peaks around $e\sim 0.8$. Beyond this enhancement we see a steep suppression which is consistent with $\tau_{e}^{-1}\propto e^{-3}$, as predicted by Papaloizou & Larwood (2000). As $e$ tends towards $0.7$ the damping rate converges towards zero. In order to better interpret the variation in these high $e$ cases, the subplots also show the variation of $\tau_{e}^{-1}$ on a logarithmic scale. This better exhibits the power law behaviour, with the predicted $\propto e^{-3}$ scaling indicated by the inset grey lines. Further connections between our present results and previous predictions will be discussed in detail in section 5.1
@@ 4.5 AMF and torque deposition function
Up to this point we have focused our attention on the torques which communicate angular momentum between the planet and the disc, causing orbital evolution. However, it should be noted that the angular momentum exchanged with the gas is not immediately deposited into the background disc. Instead the coupling is mediated by the launching of density waves, as visualized in Fig. 2, which propagate away from the planet and transport angular momentum beyond the excitation region. It is only when the waves damp via linear (Takeuchi et al., 1996; Miranda & Rafikov, 2020) or nonlinear (Goodman & Rafikov, 2001; Rafikov, 2002a; Cimerman & Rafikov, 2021) processes that they deposit angular momentum causing the disc to evolve – driving mass flow and sculpting structures such as gaps and bumps in the surface density profile (e.g. Lynden-Bell & Pringle, 1974; Rafikov, 2002b; Cordwell & Rafikov, 2024).
Our linear theory is manifestly only able to capture the linear evolution of excited density waves. For our (locally) isothermal discs a freely propagating density wave is known to be associated with a conserved wave-action such that
@@
away from the planet, at which point they start to deposit their angular momentum. This shocking behaviour is naturally captured by fully nonlinear, numerical simulations. However, one should note that simulations are also susceptible to numerical dissipation which can artificially damp the waves. At our fiducial lower resolution, damping associated with weakly nonlinear shocks was precluded by numerical dissipation , although the eccentric modification of $F_{J}$ associated with planetary excursions and the mutual interaction of waves near $r_{\rm p}$ was resolved adequately. Therefore, in order to reduce numerical dissipation, in this section we repeated a limited set of high resolution simulations with much higher radial resolution $N_{r}=18432,$ but holding the azimuthal resolution constant at $N_{\phi}\approx 6528$. In the vicinity of the planetary semi-major axis, this corresponds to $\sim 281$ cells/$H_{\rm p}$. [^4] This resolution approaches that of Cimerman & Rafikov (2021), which used 464 cells/$H_{\rm p}$ in order to achieve quantitative convergence of the shock formation. However, their planet mass was $M_{\rm p}=0.01M_{\rm th}$ and they only considered circular orbits, precluding a direct comparison.
@@
As the eccentricity is increased, the radial structure of the excited AMF notably changes. For the transonic value of $e=0.06$ we see this as a series of small ‘steps’ between the peri and apocentre locations, denoted by the brown dashed lines. Within this interval, the linear theory and simulations still agree fairly well. Once again, further out the simulation AMF decays whilst the linear theory $F_{J}$ is preserved. However, now it appears that the nonlinear damping sets in closer to the planetary semi-major axis. As we consider the fully supersonic regime with $e=0.12$, we see that the AMF about $a_{\rm p}$ actually flips to become a local maximum. This enhancement might be interpreted as the superposition of the inwards travelling waves launched from apocentre and the outwards travelling waves excited at pericentre. Beyond this, the linear theory now saturates with a higher level of $F_{J}$ in the inner disc compared with the outer disc. This inversion is consistent with the torque reversals discussed in section 4.3. Meanwhile the simulations again exhibit a clear departure from the linear theory outside of the orbital extremities. The picture is similar for $e=0.3$ where the central peak is now more spread out due to the larger radial excursions executed by the planet. Since the excitation happens over a wider range, there is also more scope for waves to shock and dissipate before propagating out of this epicylic interval. Therefore the AMF actually begins to dip beneath the linear theory predictions within the radial range of the planetary excursions as one examines close to pericentre and apocentre.
@@
which accounts for the balance between the torque density injected by the perturber and the radial variation of wave AMF. [^5] In these plots we restrict out attention to $r/a_{\rm p}>0.4$. Interior to this the interaction with the inner boundary contaminates the AMF and the behaviour is less reliable. In purely linear theory these terms cancel and there is no angular momentum deposited in the background disc. With nonlinear shocking at play, $-dF_{J}/dr$ dominates the deposition. Knowledge of the spatial distribution of $dF_{\rm dep}/{dr}$ is important for modelling the evolution of surface density, as expounded in the framework developed by Cordwell & Rafikov (2024).
@@ 5.2.1 Nonlinear corotation effects
$\tau_{\rm lib}\sim\frac{8\pi}{3x_{\rm s}}\Omega_{\rm p}^{-1}=219T_{\rm p}.$ (31)
Therefore, by sampling our simulation torques at times $T_{\rm p}<t\ll\tau_{\rm lib}$ we have aimed to probe the torques during a phase where they have had time to establish the linear response, but before saturation effects complicate proceedings.
Figure 11: AMF and torque evolution in the coorbital region for $e=0.01$. Panel (a): shows the AMF radial profile for the simulation results at different times denoted by the solid, coloured lines. The linear theory prediction is shown as the overlying dotted yellow line. Panel (b): the coorbital torque $T_{\rm C}$ contribution is plotted as a function of time. The coloured, vertical dashed lines correspond to the time slices shown in (a). Note that the y-axes here are much narrower than the typical AMF and torque scales used in previous figures, in order to highlight these subtle coorbital effects.
@@
In Fig. 11 we first visualise this time dependence for the simplest case of our near circular run with $e=0.01$, in the disc with $(p,q)=(0.5,0.0)$. In panel (a) we zoom in to the AMF distribution around the coorbital region between $|r-a_{\rm p}|<0.03$. The solid lines show the $F_{J}$ profiles measured in our simulation at different times. The colours correspond to the time slices shown in panel (b), which in turn show the time dependence of the disc integrated torque $T_{\rm C}$ within this narrow radial range [^6] Note that this integral will pick up contributions from both Lindblad and corotation resonances.. We clearly see that $T_{\rm C}$ undergoes a long time scale evolution. At $t=T_{\rm p}$ the AMF profile is rapidly evolving and the net torque profile is still settling down. About the coorbital region, $F_{J}$ is smooth at this time. However, over several dynamical timescales the AMF develops a step-like jump which is comparable to the predictions from linear theory (see the over-plotted dotted line). Within the context of linear theory, this flux ‘discontinuity’ is a measure of the corotation torque (Goldreich & Tremaine, 1979; Korycansky & Pollack, 1993) and has previously been used to partition the net torque into its Lindblad and corotation contributions (e.g. FR25). However, we should note that the torque and the AMF jump profile continue to slowly evolve, with $T_{\rm C}$ decreasing towards a minimum by $t\sim 75T_{\rm p}$. This represents the slow transition towards the unsaturated horseshoe drag regime, over a fraction of $\tau_{\rm lib}$, where there is an enhancement in the negative corotation torques, qualitatively similar to the picture presented in Fig. 7 of Paardekooper & Papaloizou (2009). At $t=100T_{\rm p}$, during this unsaturated phase, the AMF jump looks similar to the earlier times but with a slightly larger magnitude, consistent with the decrease in disc torques. Beyond this time, libration effects begin to enter as material has had time to traverse the tadpole orbits and catch up with the planet. We see that the net torque begins to rise as saturation starts setting in. Simulating for longer times would presumably show the long timescale libration oscillations similar to those seen in Fig. 12 of Paardekooper & Papaloizou (2009), before levelling off in the saturated state. We see that this phase of evolution is associated with a flattening out of the AMF and the jump is suppressed.
@@
We compare this behaviour with a run boasting the much higher eccentricity of $e=0.3$. The corresponding results are shown in Fig. 12. Once again there is an initial transient phase where the torques and AMF profiles are rapidly set up over dynamical timescales. However, by $t=25T_{\rm p}$ the $F_{J}$ discontinuity has formed but with a much smaller amplitude. Furthermore, this profile seems to undergo slower evolution compared with the near circular run discussed above. We find that the AMF curves closely overlap for all $t\geq 25T_{\rm p}$ and the $T_{\rm C}$ profile remains level towards late times. This modified picture is consistent with the view that larger eccentricities suppress the corotation torques. This has previously been seen within the context of linear theory by FR25 and also found for the nonlinear horseshoe drag regime by Fendyke & Nelson (2014). They attribute this effect to the enhanced softening scale set by the larger epicylic excursions of the planet, which in turn shrinks $x_{\rm s}$. This will increase $\tau_{\rm lib}$ and hence explains the delayed onset of saturation signatures in our more eccentric runs. Within the context of linear theory this might be interpreted as the dilution of the perturbing potential over many more, non-coorbital corotation resonances, which weakens the primary coorbital resonance. These non-coorbital corotation resonances are also predicted to exhibit small AMF jumps within the context of the linear theory (see FR25, Fig. 7) although these very subtle features are not found in our simulations. Indeed, the extended effects of nonlinear shocking lead to stronger departures from the linear theory in the region between peri and apocentre, as discussed in section 4.5. This complicates the comparison with linear AMF predictions and presumably washes out many of these subtle features.

### v2

@@ 4.3 Integrated torques
As anticipated by the agreement observed in the radial torque density curves examined in Figs. 4 and 5, here we also find close accordance in the integrated torque across all eccentricities and disc profiles. Notably, the simulations confirm the qualitative dependence on $e$ as previously discussed in detail by FR25. Namely, for the steeper $p=1.5$ runs there is an initial enhancement in $T$ as one approaches and exceeds the transonic value of $e=h_{\rm p}$. The simulations are specifically targeted to capture this peak torque around $e=0.06$, finding almost perfect overlap with the linear predictions. Meanwhile we see that for the shallower surface density profile $p=0.5$, this enhancement is less pronounced and the $T$ remains relatively flat up towards $e=h_{\rm p}$. Beyond this transonic point, we observe a clear transition in behaviour as the net torque undergoes a sharp reversal becoming highly negative and minimised when $e\sim 2h_{\rm p}$. This corresponds to the value of eccentricity for which the planet moves supersonically with respect to the gas at all orbital phases.
@@
It was at this value of $e=0.12$ that FR25 truncated their study due to uncertainty in the modal convergence for higher eccentricities. However, here we push much further and extend our simulations and linear theory up towards $e=0.7$. The net torque remains negative but slowly decreases in absolute value as $e$ increases. Fitting a power law between $e=0.18$ and $e=0.7$ to our linear theory results, yields the best fitting exponents (1.1,1.1,1.0)(-1.1,-1.1,-1.0) for the disc profiles $(p,q)=$ $(1.5,1.0)$, $(1.5,0.0)$ and $(0.5,0.0)$ respectively. This scaling is close to the heuristically motivated fitting functions presented in Ida et al. (2020), which suggest that $T\propto e^{-1}$ in the highly supersonic limit. Once again, the broad trends are in good agreement between the methodologies. However, for the highest values of $e$ we begin to see larger discrepancies emerge as the linear theory underestimates the torque magnitude found in the simulations. To test this further, we have performed a convergence study in appendix A which demonstrates that our linear framework and simulation torque diagnostics do not change significantly upon including more modes or doubling the fiducial resolution. This suggests that the differences arise due to more intrinsic difficulties. Perhaps the high value of $e$ fundamentally precludes convergence of our modal scheme. Furthermore, the very oscillatory nature of the wave contributions for high order modes becomes mutually challenging for the linear theory ODE solvers and the fully numerical hydro scheme. Whilst the large radial planetary excursions potentially lead to more troublesome interaction with the boundaries, we show in appendix A.3 that the differences are primarily sourced between apocentre and pericentre. The intricate torque structures in this interval might point towards subtle nonlinear effects that affect the disc response in this region. For example, in section 4.5 we show that nonlinear wave steepening can deposit angular momentum into the disc, which will modify the background density structure slightly, potentially altering the interaction with the planet.
@@ 4.4 Orbital evolution
In Figs. 8 and 9 we examine this orbital element interplay for $\tau_{a}^{-1}$ and $\tau_{e}^{-1}$ respectively, as a function of $e$. Once again, the three rows in each figure correspond to the different disc profiles. Meanwhile the yellow markers and connective lines correspond to the linear theory predictions, whilst the blue stars mark the simulation results. The results are plotted in units of the characteristic migration rate $\tau_{0}^{-1}\equiv F_{J,0}/(M_{\rm p}n_{\rm p}a_{\rm p}^{2})$.
@@
Now turning our attention to Fig. 9, we examine the behaviour of the characteristic eccentricity damping rate $\tau_{e}^{-1}$. Firstly, note the different y-axis scale which is roughly two orders of magnitudes larger compared with Fig. 8, consistent with our expectation that the $\tau_{e}^{-1}$ is much faster than $\tau_{a}^{-1}$ (as found previously by Papaloizou & Larwood, 2000; Tanaka & Ward, 2004). Once again we see the simulations and linear results are in very good agreement. The characteristic eccentricity damping exhibits a sharp rise which peaks around $e\sim 0.08$. Beyond this enhancement we see a steep suppression which is consistent with $\tau_{e}^{-1}\propto e^{-3}$, as predicted by Papaloizou & Larwood (2000). As $e$ tends towards $0.7$ the damping rate converges towards zero. In order to better interpret the variation in these high $e$ cases, the subplots also show the variation of $\tau_{e}^{-1}$ on a logarithmic scale. This better exhibits the power law behaviour, with the predicted $\propto e^{-3}$ scaling indicated by the inset grey lines. Further connections between our present results and previous predictions will be discussed in detail in section 5.1
@@ 4.5 AMF and torque deposition function
Up to this point we have focused our attention on the torques which communicate angular momentum between the planet and the disc, causing orbital evolution. However, it should be noted that the angular momentum exchanged with the gas is not immediately deposited into the background disc. Instead the coupling is mediated by the launching of density waves, as visualized in Fig. 2, which propagate away from the planet and transport angular momentum beyond the excitation region. It is only when the waves damp via linear (Takeuchi et al., 1996; Miranda & Rafikov, 2020) or nonlinear (Goodman & Rafikov, 2001; Rafikov, 2002a; Cimerman & Rafikov, 2021) processes that they deposit angular momentum causing the disc to evolve – driving mass flow and sculpting structures such as gaps and bumps in the surface density profile (e.g. Lynden-Bell & Pringle, 1974; Rafikov, 2002b; Dempsey et al., 2020; Cordwell & Rafikov, 2024).
Our linear theory is manifestly only able to capture the linear evolution of excited density waves. For our (locally) isothermal discs a freely propagating density wave is associated with a conserved wave-action (e.g. Lee, 2016; Miranda & Rafikov, 2020) such that
@@
[…]
@@
As the eccentricity is increased, the radial structure of the excited AMF notably changes. For the transonic value of $e=0.06$ we see this as a series of small ‘steps’ between the peri and apocentre locations, denoted by the brown dashed lines. Within this interval, the linear theory and simulations still agree fairly well. Once again, further out the simulation AMF decays whilst the linear theory predicts that $F_{J}$ is preserved. However, now it appears that the nonlinear damping sets in closer to the planetary semi-major axis. As we consider the fully supersonic regime with $e=0.12$, we see that the AMF about $a_{\rm p}$ actually flips to become a local maximum. This enhancement might be interpreted as the superposition of the inwards travelling waves launched from apocentre and the outwards travelling waves excited at pericentre. Beyond this, the linear theory now saturates with a higher level of $F_{J}$ in the inner disc compared with the outer disc. This inversion is consistent with the torque reversals discussed in section 4.3. Meanwhile the simulations again exhibit a clear departure from the linear theory outside of the orbital extremities. The picture is similar for $e=0.3$ where the central peak is now more spread out due to the larger radial excursions executed by the planet. Since the excitation happens over a wider range, there is also more scope for waves to shock and dissipate before propagating out of this epicylic interval. Therefore the AMF actually begins to dip beneath the linear theory predictions within the radial range of the planetary excursions as one examines close to pericentre and apocentre.
@@
which accounts for the balance between the torque density injected by the perturber and the radial variation of wave AMF. [^6] In these plots we restrict out attention to $r/a_{\rm p}>0.4$. Interior to this the interaction with the inner boundary contaminates the AMF and the behaviour is less reliable. In purely linear theory these terms cancel and there is no angular momentum deposited in the background disc. With nonlinear shocking at play, $dF_{J}/dr$ dominates the deposition. Knowledge of the spatial distribution of $dF_{\rm dep}/{dr}$ is important for modelling the evolution of surface density, as expounded in the framework developed by Cordwell & Rafikov (2024).
@@ 5.2.1 Nonlinear corotation effects
$\tau_{\rm lib}\sim\frac{8\pi}{3x_{\rm s}}n_{\rm p}^{-1}\simeq 220T_{\rm p}.$ (31)
Therefore, up until now, by sampling our simulation torques at times $t$ satisfying $T_{\rm p}<t\ll\tau_{\rm lib}$ we have aimed to probe the torques during a phase where they have had time to establish the linear response, but before saturation effects complicate proceedings. However, we now run a selection of extended runs up towards 200 planetary orbits in order to start probing these late time, nonlinear effects.
Figure 12: AMF and torque evolution in the coorbital region for $e=0.01$. Panel (a): shows the AMF radial profile for the simulation results at different times denoted by the solid, coloured lines. The linear theory prediction is shown as the overlying dotted yellow line. Panel (b): the coorbital torque $T_{r\sim a_{\rm p}}$ contribution is plotted as a function of time. The coloured, vertical dashed lines correspond to the time slices shown in (a). Note that the y-axes here are much narrower than the typical AMF and torque scales used in previous figures, in order to highlight these subtle coorbital effects.
@@
In Fig. 12 we first visualise this time dependence for the simplest case of our near circular run with $e=0.01$, in the disc with $(p,q)=(0.5,0.0)$. In panel (a) we zoom in to the AMF distribution around the coorbital region between $|r-a_{\rm p}|<0.03$. The solid lines show the $F_{J}$ profiles measured in our simulation at different times. The colours correspond to the time slices shown in panel (b), which in turn show the time dependence of the disc integrated torque $T_{r\sim a_{\rm p}}$ within this narrow radial range [^7] Note that this integral will pick up contributions from both Lindblad and corotation resonances.. We clearly see that $T_{r\sim a_{\rm p}}$ undergoes a long time scale evolution. At $t=T_{\rm p}$ the AMF profile is rapidly evolving and the net torque profile is still settling down. About the coorbital region, $F_{J}$ is smooth at this time. However, over several dynamical timescales the AMF develops a step-like jump which is comparable to the predictions from linear theory (see the over-plotted dotted line). Within the context of linear theory, this flux ‘discontinuity’ is a measure of the corotation torque (Goldreich & Tremaine, 1979; Korycansky & Pollack, 1993) and has previously been used to partition the net torque into its Lindblad and corotation contributions (e.g. FR25). However, we should note that the torque and the AMF jump profile continue to slowly evolve, with $T_{r\sim a_{\rm p}}$ decreasing towards a minimum by $t\sim 75T_{\rm p}$. This represents the slow transition towards the unsaturated horseshoe drag regime, over a fraction of $\tau_{\rm lib}$, where there is an enhancement in the negative corotation torques, qualitatively similar to the picture presented in Fig. 7 of Paardekooper & Papaloizou (2009). At $t=100T_{\rm p}$, during this unsaturated phase, the AMF jump looks similar to the earlier times but with a slightly larger magnitude, consistent with the decrease in disc torques. Beyond this time, libration effects begin to enter as material has had time to traverse the tadpole orbits and catch up with the planet. We see that the net torque begins to rise as saturation starts setting in. Simulating for longer times would presumably show the long timescale libration oscillations similar to those seen in Fig. 12 of Paardekooper & Papaloizou (2009), before levelling off in the saturated state. We see that this phase of evolution is associated with a flattening out of the AMF and the jump is suppressed.
@@
We compare this behaviour with a run boasting the much higher eccentricity of $e=0.3$. The corresponding results are shown in Fig. 13. Once again there is an initial transient phase where the torques and AMF profiles are rapidly set up over dynamical timescales. However, by $t=25T_{\rm p}$ the $F_{J}$ discontinuity has formed but with a much smaller amplitude. Furthermore, this profile seems to undergo slower evolution compared with the near circular run discussed above. We find that the AMF curves closely overlap for all $t\geq 25T_{\rm p}$ and the $T_{r\sim a_{\rm p}}$ profile remains level towards late times. This modified picture is consistent with the view that larger eccentricities suppress the corotation torques. This has previously been seen within the context of linear theory by FR25 and also found for the nonlinear horseshoe drag regime by Fendyke & Nelson (2014). They attribute this effect to the enhanced softening scale set by the larger epicylic excursions of the planet, which in turn shrinks $x_{\rm s}$. This will increase $\tau_{\rm lib}$ and hence explains the delayed onset of saturation signatures in our more eccentric runs. Within the context of linear theory this might be interpreted as the dilution of the perturbing potential over many more, non-coorbital corotation resonances, which weakens the primary coorbital resonance. These non-coorbital corotation resonances are also predicted to exhibit small AMF jumps within the context of the linear theory (see FR25, Fig. 7) although these very subtle features are not found in our simulations. Indeed, the extended effects of nonlinear shocking lead to stronger departures from the linear theory in the region between peri and apocentre, as discussed in section 4.5. This complicates the comparison with linear AMF predictions and presumably washes out many of these subtle features.